A 5 kg block of 20cm × 10 cm × 8 cm dimensionless lying on the largest base . It is now made to stand with length vertical . If g=10. Then the amount of work done is

Grade:11

1 Answers

Rohana
6 years ago
Work done = f*d 
F = m*a. 
F = 5*10=50N.  
Work done = 50*0.2 = 10J.  
Thank you I think it is correct please approve the answer